# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ 3.1 Theoretical Guarantee
This subsection shows that the value of the expected discounted cumulative reward produced by ${\pi}^{\mathrm{est}}_{k,m}$ is approximately optimal, where the optimality gap decays as $k\!\to\!n$ and $m$ becomes large.
@@
Bellman noise. We introduce the notion of Bellman noise, which is used in the main theorem. Consider $\widehat{\mathcal{T}}_{k,m}$. Clearly, it is an unbiased estimator of the generalized adapted Bellman operator $\widehat{\mathcal{T}}_{k}$,
@@
$\begin{split}&\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})\!=\!r_{\Delta}(s,a)\!+\!\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in\Delta\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime}).\end{split}$ (16)
@@
For all $(s_{g},s_{\Delta},a_{g},a_{\Delta})\!\in\!\mathcal{S}_{g}\times\!\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$, set $\widehat{Q}^{0}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0$. For $t\in\mathbb{N}$, let $\widehat{Q}_{k}^{t+1}=\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{t}$, where $\widehat{\mathcal{T}}_{k}$ is defined for $k\leq n$ in Eq. 16. Then, $\widehat{\mathcal{T}}_{k}$ is also a $\gamma$-contraction (Lemma B.7) with fixed-point $\widehat{Q}_{k}^{*}$. By the law of large numbers, $\lim_{m\to\infty}\widehat{\mathcal{T}}_{k,m}\!=\!\widehat{\mathcal{T}}_{k}$ and $\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}\to 0$ as $m\to\infty$. For finite $m$, $\epsilon_{k,m}\coloneq\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}$ is the well-studied Bellman noise.
Lemma 3.2.
By the Chernoff bound, for $k\in[n]$ and $m\in\mathbb{N}$, where $m$ is the number of samples in Eq. 10, $\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}\leq\epsilon_{k,m}\leq\widetilde{O}(1/\sqrt{k})$, when $m=m^{*}=\frac{2|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{2.5+|\mathcal{S}_{l}||\mathcal{A}_{l}|}}{(1-\gamma)^{5}}\log(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$.
We defer the proof of Lemma 3.2 to Appendix F.1. Next, let ${\pi}_{k}^{\mathrm{est}}\coloneq\widetilde{\pi}_{k,m^{*}}^{\mathrm{est}}$. To compare the difference in the performance of $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$, we define the value function of a policy $\pi$ by $V^{\pi}$:
@@
$V^{\pi}(s)=\mathbb{E}_{\begin{subarray}{c}a(t)\sim\pi(\cdot|s(t))\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))\bigg{|}s(0)=s\right].$ (17)
Intuitively, the value function $V^{\pi}(s)$ is the expected discounted cumulative reward when starting from state $s$ and applying actions from the policy $\pi$ across an infinite horizon. With the above preparations, we are primed to present our main result: a decaying bound on the optimality gap for our learned policy ${\pi}_{k}^{\mathrm{est}}$.
@@
Let $\widetilde{\pi}_{k}$ denote the learned policy from SUBSAMPLE-MFQ. Then, $\forall s\in\mathcal{S}$:
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k}^{\mathrm{est}}}(s_{0})\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\widetilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{21}{10\sqrt{k}}\leq\widetilde{O}(1/\sqrt{k})$
We provide for the theorem in Appendix C, and defer the proof to Appendix F.
Discussion 3.5.
Between Algorithms 2 and 1, the asymptotic sample complexity to learn $\widehat{\pi}^{\mathrm{est}}_{k}$ for a fixed $k$ is $\min\{O(|\mathcal{Z}_{l}|^{k}),O(k^{|\mathcal{Z}_{l}|})\}$. By Theorem 3.4, as $k\to n$, the optimality gap decays, revealing a fundamental trade-off in the choice of $k$: increasing $k$ improves the performance of the policy, but increases the size of the $Q$-function. We explore this trade-off further in our experiments. For $k=O(\log n)$, the runtime is $\min\{O(n^{\log|\mathcal{Z}_{l}|}),O((\log n)^{|\mathcal{Z}_{l}|})\}$. This is an exponential speedup on the complexity from mean-field value iteration (from $\mathrm{poly}(n)$ to $\mathrm{poly}(\log n)$), as well as over traditional value-iteration (from $\mathrm{exp}(n)$ to $\mathrm{poly}(n)$). Further, the optimality gap decays to 00 at the rate of $O(1/\sqrt{\log n})$.
Appendix G extends the theorem to stochastic reward distributions.
@@
Suppose we are given two families of distributions on the reward functions, $r_{g}(s_{g},a_{g})\sim\{\mathcal{G}_{s_{g},a_{g}}\}_{s_{g},a_{g}\in\mathcal{S}_{g}\times\mathcal{A}_{g}}$ and $r_{l}(s_{i},s_{g},a_{i})\sim\{\mathcal{L}_{s_{i},s_{g},a_{i}}\}_{s_{i},s_{g},a_{i}\in\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}}$. Then, under standard assumptions of boundedness of the support of $\mathcal{G}_{s_{g},a_{g}}$ and $\mathcal{L}_{s_{i},s_{g},a_{i}}$, SUBSAMPLE-MFQ learns a stochastic policy $\pi_{k}^{\mathrm{est}}$ satisfying
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\pi^{\mathrm{est}}_{k}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}}.$
@@
In the non-tabular setting with infinite state/action spaces, one could replace the $Q$-learning algorithm with any arbitrary value-based RL method that learns $\widehat{Q}_{k}$ with function approximation (Sutton et al., 1999) such as deep $Q$-networks (Silver et al., 2016). Doing so raises an additional error that factors into Theorem 3.4.
Definition 3.7 (Linear MDP).
@@
$\mathrm{MDP}(S,A,\mathbb{P},r)$ is a linear MDP with feature map $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ if there exist $d$ unknown (signed) measures $\mu=(\mu^{1},\dots,\mu^{d})$ over $\mathcal{S}$ and a vector $\theta\in\mathbb{R}^{d}$ such that for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, we have $\mathbb{P}(\cdot|s,a)=\langle\phi(s,a),\mu(\cdot)\rangle$ and $r(s,a)=\langle\phi(s,a),\theta\rangle$.
@@
Suppose the system is a linear MDP, where $\mathcal{S}_{g}$ and $\mathcal{S}_{l}$ are infinite compact sets. By a reduction from Ren et al. (2024) and using function approximation to learn the spectral features $\phi_{k}$ for $\widehat{Q}_{k}$, we derive a performance guarantee for the learned policy $\pi_{k}^{\mathrm{est}}$, where the optimality gap decays with $k$.
@@
When ${\pi}^{\mathrm{est}}_{k}$ is derived from the spectral features $\phi_{k}$ learned in $\widehat{Q}_{k}$, and $M$ is the number of samples used in the linear function approximation,
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}^{\mathrm{est}}_{k}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}+\log\left(2k^{2}\right)\frac{\|{\phi}_{k}\|^{5}}{\sqrt{M}}+\frac{2}{\sqrt{k}}\cdot\frac{\gamma\widetilde{r}}{1-\gamma}\|\bar{\phi}_{k}\|\right)\right]\geq\left(1-\frac{1}{100\sqrt{k}}\right)\cdot\left(1-\frac{2}{\sqrt{k}}\right)$
We defer the proof of Theorem 3.8 to Appendix H.
Remark 3.9.
If $k=O(\log n)$, SUBSAMPLE-MFQ can handle $|\mathcal{E}|=O(\log n/\log\log n)$ different types of local agents, since the run-time of the learning algorithm becomes $\mathrm{poly}(n)$. This additionally supersedes the previous-best heterogeneity capacity from Mondal et al. (2022), which only handles constant $|\mathcal{E}|$.
Remark 3.10.
Our algorithm and policy also contributes to the growing literature on the centralized-training-decentralized-execution paradigm (Zhou et al., 2023; Wang et al., 2023), and the literature on exogenous MDPs, wherein our algorithm has an advantage in that the agents do not all have to “see” each other during the online (execution) stage, and hence contains a partially observable setting.
@@ 4 Conclusion and Future Works
This work develops subsampling for mean field MARL in a cooperative system with a global decision-making agent and $n$ homogeneous local agents. We propose SUBSAMPLE-MFQ which learns each agent’s best response to the mean effect from a sample of its neighbors, allowing an exponential reduction on the sample complexity of approximating a solution to the MDP. We provide a theoretical analysis on the optimality gap of the learned policy, showing that the learned policy converges to the optimal policy with the number of agents $k$ sampled at the rate $\widetilde{O}(1/\sqrt{k})$ validate our theoretical results through numerical experiments. We further extend this result to the non-tabular setting with infinite state and action spaces.
We recognize several future directions. Firstly, this model studies a ‘star-network’ setting to model a single source of density. It would be fascinating to extend this subsampling framework to general networks. We believe expander-graph decompositions (Anand & Umans, 2023; Reingold, 2008) are amenable for this. A second direction would be to find connections between our sub-sampling method to algorithms in federated learning, where the rewards can be stochastic. A third direction of this work would be to consider the setting of truly heterogeneous local agents. Finally, it would be exciting to generalize this work to the online setting without a generative oracle: we conjecture that tools from recent works on stochastic approximation (Chen & Theja Maguluri, 2022) and no-regret RL (Jin et al., 2021b) might be valuable.

### v3

@@ 4 Theoretical Guarantees and Analysis Approach
We now show the value of the expected discounted cumulative reward produced by ${\pi}^{\mathrm{est}}_{k,m}$ is approximately optimal, where the optimality gap decays as $k\to n$ and $m$ grows.
@@
Bellman noise. We introduce the notion of Bellman noise, which is used in the main theorem. Note that $\hat{\mathcal{T}}_{k,m}$ is an unbiased estimator of the adapted Bellman operator $\hat{\mathcal{T}}_{k}$,
@@
$\begin{split}&\hat{\mathcal{T}}_{k}\hat{Q}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})\!=\!r_{\Delta}(s,a)\!+\!\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot\mid s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot\mid s_{i},s_{g},a_{i}),\forall i\in\Delta\end{subarray}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime}).\end{split}$ (11)
@@
Let $\smash{\hat{Q}^{0}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0}$. For $\smash{t>0}$, let $\smash{\hat{Q}_{k}^{t+1}=\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{t}}$, where $\hat{\mathcal{T}}_{k}$ is defined for $k\leq n$ in Equation 11. Then, $\hat{\mathcal{T}}_{k}$ is also a $\gamma$-contraction with fixed-point $\hat{Q}_{k}^{*}$. By the law of large numbers, $\smash{\lim_{m\to\infty}\hat{\mathcal{T}}_{k,m}=\hat{\mathcal{T}}_{k}}$ and $\smash{\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}\to 0}$ as $m\to\infty$. For finite $m$, $\epsilon_{k,m}\coloneq\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}$ is the well-studied Bellman noise. To compare the performance between $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$, we define the value function of a policy $\pi$:
@@
For a given policy $\pi$, the value function $V^{\pi}:\mathcal{S}\to\mathbb{R}$ for $\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ is given by:
$V^{\pi}(s)\!=\!\mathop{\mathbb{E}}_{\begin{subarray}{c}a(t)\sim\pi(\cdot|s(t))\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))\bigg{|}s(0)\!=\!s\right]\!\!.\!$ (12)
Intuitively, $V^{\pi}(s)$ is the expected discounted cumulative reward when starting from state $s$ and applying actions from the policy $\pi$ across an infinite horizon.
With the above preparations, we are primed to present our main result: a high-probability bound on the optimality gap for our learned policy ${\pi}_{k,m}^{\mathrm{est}}$ that decays with rate $\tilde{O}(1/\sqrt{k})$.
@@
Let $\pi_{k,m}^{\mathrm{est}}$ denote the learned policy deployed in SUBSAMPLE-MFQ: Execution. Then, for all $s_{0}\in\mathcal{S}\coloneqq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, we have
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}^{\mathrm{est}}}(s_{0})\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\tilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{1}{10\sqrt{k}}+2\epsilon_{k,m}$
We prove Theorem 4.4 in Appendix G, and provide a proof sketch in Appendix D. We also generalize the result to stochastic rewards in Appendix H.
To control the Bellman noise $\epsilon_{k,m}$, we show that for sufficiently many samples $m^{*}$, $\epsilon_{k,m^{*}}$ also decays on the order of $\tilde{O}(1/\sqrt{k})$ with high probability. For this, we introduce Lemma 4.3:
Lemma 4.3 (Controlling the Bellman Noise.).
For $k\in[n]$, let
$m^{*}=2|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{3.5+|\mathcal{S}_{l}||\mathcal{A}_{l}|}\frac{\log({|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|})}{(1-\gamma)^{5}}\log\frac{1}{(1-\gamma)^{2}}$
be the number of samples in Equation 8. If the number of iterations $T$ satisfies $\smash{T\geq\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}}$ then with probability at least $\smash{1-\frac{1}{100e^{k}}}$, the Bellman error satisfies $\smash{\epsilon_{k,m^{*}}\leq\tilde{O}(\frac{1}{\sqrt{k}})}$.
We defer the proof of Lemma 4.3 to Section G.1. To simplify notation, let ${\pi}_{k}^{\mathrm{est}}\coloneq{\pi}_{k,m^{*}}^{\mathrm{est}}$. Then, by combining Lemma 4.3 and Theorem 4.2, we obtain our main result in Theorem 4.4:
@@
Let ${\pi}^{\mathrm{est}}_{k}$ denote the learned policy from SUBSAMPLE-MFQ: Execution where the number of samples $m$ is determined in Lemma 4.3. Suppose $T\geq\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$. Then, $\forall s_{0}\in\mathcal{S}\coloneq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, with probability at least $1-1/100e^{k}$, we have  [^1] The $1/100e^{k}$ term can be replaced to any arbitrary $\delta>0$ at the cost of attaching logarithmic dependencies of $\delta$ on the $4/\sqrt{k}$ term in the error bound.,
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k}^{\mathrm{est}}}(s_{0})\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}\ln\frac{40\tilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}\!+\!\frac{4}{\sqrt{k}}.$
Our analysis hinges on two non-trivial technical steps. First, for an intermediary step in Theorem E.3 we establish that TV distance, rather than the stronger KL-divergence, is the correct metric to use (as the KL-divergence between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$ decays too slowly as $k\to n$, as we show in Lemma F.6). This requires exploiting a recent extension of the DKW inequality to sampling without replacement (Anand and Qu, 2024), and showing that the transition dynamics we study satisfies the data-processing inequality for TV-distance [^2] This is not true for general transition functions, since TV-distance is not an $f$-divergence.. Secondly, we adapt the celebrated performance-difference lemma (Kakade and Langford, 2002) to our multi-agent setting, which entails a principled analysis and careful probabilistic argument (to which we refer the reader to Appendix G).
Remark 4.5.
We also extend the formulation of SUBSAMPLE-MFQ to off-policy $Q$-learning Chen et al. (2021b); Chen and Maguluri (2022a), which replaces the generative oracle assumption with a stochastic approximation scheme that learns an approximately optimal policy using historical data. Appendix I provides theoretical guarantees with a similar decaying optimality gap as in Theorem 4.4.
Remark 4.6.
The asymptotic sample complexity of Algorithm 1 for learning $\hat{\pi}^{\mathrm{est}}_{k}$ for a fixed $k$ is $\min\{\tilde{O}(|\mathcal{Z}_{l}|^{k}|\mathcal{S}_{g}||\mathcal{A}_{g}|),\tilde{O}(k^{|\mathcal{Z}_{l}|}|\mathcal{Z}_{l}||\mathcal{S}_{g}||\mathcal{A}_{g}|)\}$, which is at least polynomially faster the standard $Q$-learning or mean-field value iteration as discussed in Remark 3.1. By Theorem 4.4, as $k\to n$, the optimality gap decays, revealing a fundamental trade-off in the choice of $k$: increasing $k$ improves the performance of the policy, but increases the size of the $Q$-function. We explore this trade-off further in our experiments. If we set $k=O(\log n)$, this leads to a sample complexity of $\min\{\tilde{O}(n^{\log|\mathcal{Z}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|),\tilde{O}((\log n)^{|\mathcal{Z}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|)\}$. This is an exponential speedup on the complexity from mean-field value iteration (from $\mathrm{poly}(n)$ to $\mathrm{poly}(\log n)$), as well as over traditional value-iteration (from $\mathrm{exp}(n)$ to $\mathrm{poly}(n)$), where the optimality gap decays to 00 with rate $O(\frac{1}{\sqrt{\log n}})$.
Remark 4.7.
If $k=O(\log n)$, SUBSAMPLE-MFQ handles $|\mathcal{E}|\leq O(\log n/\log\log n)$ types of local agents, since the run-time of the learning algorithm becomes $\mathrm{poly}(n)$. This surpasses the previous heterogeneity capacity from Mondal et al. (2022), which only handles constant $|\mathcal{E}|\leq O(1)$.
@@
In the non-tabular setting with infinite state/action spaces, one could replace the $Q$-learning algorithm with any arbitrary value-based RL method that learns $\hat{Q}_{k}$ with function approximation (Sutton et al., 1999) such as deep $Q$-networks (Silver et al., 2016). Doing so raises an additional error that factors into Theorem 4.4. We formalize this below.
Assumption 4.8 (Linear MDP with infinite state spaces).
@@
Suppose $\mathcal{S}_{g}$ and $\mathcal{S}_{l}$ are infinite compact sets. Furthermore, suppose there exists a feature map $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ and $d$ unknown (signed) measures $\mu=(\mu^{1},\dots,\mu^{d})$ over $\mathcal{S}$ and a vector $\theta\in\mathbb{R}^{d}$ such that for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, we have $\mathbb{P}(\cdot|s,a)=\langle\phi(s,a),\mu(\cdot)\rangle$ and $r(s,a)=\langle\phi(s,a),\theta\rangle$.
The existence of $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ implies one can estimate the $Q$-function of any policy as a linear function. This assumption is commonly used in policy iteration algorithms Lattimore et al. (2020); Wang et al. (2023), and allows one to obtain sample complexity bounds that are independent of $|\mathcal{S}_{l}|$ and $|\mathcal{A}_{l}|$. Finally, as is standard in RL, we assume bounded feature-norms (Tkachuk et al., 2023):
Assumption 4.9 (Bounded features).
We assume that $\|\phi(s,a)\|_{2}\leq 1$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$.
@@
Then, through a reduction from Zhang et al. (2024); Ren et al. (2024) that uses function approximation to learn the spectral features $\phi_{k}$ for $\hat{Q}_{k}$, we derive a performance guarantee for the learned policy $\pi_{k}^{\mathrm{est}}$, where the optimality gap decays with $k$.
@@
When ${\pi}^{\mathrm{est}}_{k}$ is derived from the spectral features $\phi_{k}$ learned in $\hat{Q}_{k}$, and $M$ is the number of samples used in the function approximation, then
$\Pr\bigg{[}V^{\pi^{*}}(s)-V^{{\pi}^{\mathrm{est}}_{k}}(s)\leq\tilde{O}\bigg{(}\frac{1}{\sqrt{k}}+\frac{\|{\phi}_{k}\|^{5}\log 2k^{2}}{\sqrt{M}}+\frac{2\gamma\tilde{r}\cdot\|\phi_{k}\|}{(1-\gamma)\sqrt{k}}\bigg{)}\bigg{]}\geq 1+\frac{200}{k}-\frac{201}{200\sqrt{k}}$
We defer the proof of Theorem 4.10 to Appendix J.
@@ 5 Conclusion and Future Works
This work develops subsampling for mean field MARL in a cooperative system with a global decision-making agent and $n$ homogeneous local agents. We propose SUBSAMPLE-MFQ which learns each agent’s best response to the mean effect from a sample of its neighbors, allowing an exponential reduction on the sample complexity of approximating a solution to the MDP. We provide a theoretical analysis on the optimality gap of the learned policy, showing that (with high probability) the learned policy converges to the optimal policy with the number of agents $k$ sampled at the rate $\tilde{O}(1/\sqrt{k})$, and validate our theoretical results through numerical experiments. We show that the decay rate is maintained, on expectation, when the reward functions are stochastic and when agents learn from a single trajectory on historical data through off-policy $Q$-learning. Finally, we extend this result to the non-tabular setting with infinite state and action spaces under assumptions of a linear MDP model.
Limitations and future work.
Our current work assumes that the global and local agents cooperate to optimize a structured reward under a specific dynamic model. While this model is more general than the federated learning setting, one direction would be to extend our algorithms and analysis to weaker network assumptions. We believe our framework, which can handle dense subgraphs, as well as expander-graph decompositions (Anand and Umans, 2023) may be amenable for this. Secondly, our current work incorporates mild heterogeneity among agents and assumes they are cooperative; thus, another avenue would be to consider settings with competitive agents or more complex agent heterogeneity. Finally, it would be exciting to generalize this work to the online no-regret setting.
Societal Impacts.
This work is theoretical and foundational in nature. As such, while it enables more scalable multi-agent algorithms, it is not tied to any specific applications or deployments.